A rectangular carbon block has dimensions 2.0 x 2.0 x 25 cm. What is the resistance between the square ends at 60 degrees C? The resistivity of carbon at 20 degrees C is 35 mW . m and the temperature coefficient of resistivity at 20 degrees C is -0.0005 per degree C.
Answer #1

R = pL/A

where p = resistivity

L = length , A = area

R = 0.035*0.25 / 0.02*0.02

= 21.875 ohm

R = Ro(1+ T)

R = 21.875(1 + (-0.0005)(60 - 20))

R = 21.4375 ohm

so the resistance between the square ends at 60 oC is 21.4375 ohm